OpenStudy (anonymous):

Find the volume of the sphere shown. Give each answer rounded to the nearest cubic unit. *sphere with radius of 5* HEELLPPP

OpenStudy (mathstudent55):

\( V_{sphere} = \dfrac{4}{3} \pi r^3\)

OpenStudy (anonymous):

so is the answer 524 ?

OpenStudy (zale101):

V= (4/3)(pi)(radius)^3 V=(4/3)(pi)(5)^3 pi=3.14

OpenStudy (zale101):

523.3 yup, you're right

OpenStudy (anonymous):

thanks

OpenStudy (zale101):

no problem